# Substrate Scope Contrastive Learning: Repurposing Human Bias to Learn Atomic Representations

### v1

@@ 4 Discussion
We have demonstrated that substrate scope tables from journal articles, known for being small and biased towards high-yielding examples, are in fact valuable sources of information for molecular representation learning. This is the first approach to utilize these groupings as an information source and to demonstrate their correlation with reactivity features and expert intuition. By learning patterns of human bias, ContraScope extracts the expert knowledge and reactivity trends underlying this bias. We illustrate this both qualitatively and quantitatively, showing how our learned embeddings align with established physical organic chemistry descriptors.
@@
While the applications showcased in this paper—such as the supervised learning of reaction yields and regioselectivity, along with the diverse substrate selection for experimental design—underscore the versatility of our learned embeddings, it is important to recognize that they serve as proofs-of-concept; our embedding approach, like any other, is not a universal solution. For instance, a nearest neighbor regression analysis of yield across all training substrate scopes indicates that two molecules with similar embeddings do not necessarily exhibit similar yields, which is similar to other common features (Figure S14). Moreover, substrates may achieve low reaction yields as a result of side reactions rather than inherent lack of reactivity at the C–X motif, which would not be captured in our approach; this may partially explain the observed lack of correlation for structurally complex aryl halides in the chemistry informer library 61 (Figure S15). Additionally, our operational presumption that all unobserved molecules in scope tables are less similar is a simplification that holds true only on average. From a practical standpoint, our loss function combines a term for anchor-positive loss with one for anchor-negative loss in a linear manner, which poses challenges in training both loss terms simultaneously and can at times lead to instability in the training process. All those challenges underscore that understanding and predicting chemical reactivity remains an open question, highlighting the need for further refinement of pre-training strategies and loss functions, specially designed for chemical data.
Overall, our work presents a conceptual framework for pre-training reactivity-aligned atomic representations from the selective bias inherent in substrate scope tables. With the availability of our code and pre-trained model, this methodology can be applied to new reaction systems and to develop novel chemical embeddings for diverse molecular classes. We are optimistic that our framework will serve as a catalyst for further advancements in the field, fostering more targeted and functionality-aligned approaches in molecular representation learning.
@@ 5 Methods
Substrate data for training In this study, we focus on aryl halides. We partitioned reactions recorded in the CAS Content Collection between 2010 and 2015 into substrate scope tables comprised of reactions from the same publication source, wherein the only variable is a single aryl halide substrate. All other substrates and recorded conditions remained constant within each scope. Importantly, all reactions take place at an aryl C–X bond. Scopes with fewer than 5 reactions and reactions without recorded yields were excluded. This led to our final training dataset of 20,798 distinct aryl halides, covering 64,192 reactions and 6,919 substrate scopes.
Calculation of reactivity descriptors All reactivity descriptors are calculated by density-functional theory (DFT) with Gaussian 1656 via an automated descriptor generation pipeline built on top of AutoQChem57. For each unique aryl halide, up to 20 conformers were generated using RDKit’s ETKDG algorithm63 and optimized using the MMFF94 force field64. To reduce computational overhead when using DFT, the lowest-energy conformer for each molecule was then selected via GFN2-xTB65. Any conformer for which any energy calculation did not converge was discarded. The lowest-energy conformer for each aryl halide then underwent geometry optimization and frequency calculations in Gaussian 16 using the B3LYP functional66, 67 with the 6-31G*68 basis set. Atoms with atomic number $>35$ use the LANL2DZ69 basis set instead. This led to the generation of 25 molecule-level descriptors per molecule (energies, energy corrections, dipole moment, HOMO/LUMO energies, electronegativity, etc.) and 19 atom-level descriptors per atom per molecule (buried volume, partial charges, NMR shielding constants, etc.), and the relevant atom-level descriptors for the reactive C–X bond were extracted. The full set of descriptors is listed in Figures S5 and S6.
Featurization and network architecture We adopt a graph representation of molecules and the graph isomorphism network (GIN)48 for modeling them. GIN is one kind of graph neural network that updates the representation of each atom over multiple iterations as follows:
@@
where $h_{v}^{(k)}$ is the atom representation of atom $v$ at $k$-th layer, $\mathcal{N}(v)$ is a set of atoms that are covalently bonded to atom $v$. $\epsilon$ is a learnable scalar and MLP represents a multi-layer preceptrons model. Utilizing this atom-wise message-passing schema, GIN is able to propagate and aggregate information through network layers to embed atoms in a manner that respects graph isomorphism. After hyper-parameter tuning, the network has a hidden dimension of 64, and message passing layer of 5. For the initial representation of molecular graphs, we use minimal featurizations49 that include atomic numbers, total degrees, formal charges, chiral tags, number of hydrogens, hybridization types, aromaticity, mass, and bond types.
Substrate scope contrastive learning The training of substrate scope contrastive learning involves triplet sampling, computing loss function depicted in Eq. 1, and back-propagation. Within an epoch, each aryl halide serves as the anchor molecule once and 16 triplets are sampled randomly for each anchor. We exclude all cases with identical molecules sampled in one triplet. Based on empirical performance, we adopt a distance metric called signal-noise ratio distance (SNR) 70, defined as follows:
@@
where ${f}_{i},{f}_{j}$ is the embedding of aryl halide $m_{i},m_{j}$. We take the representation of an aryl halide’s reactivity, $f_{i}$, to be the atom-level feature of the carbon atom in the reactive C–X bond from the GIN. The Adam71 optimization algorithm is used for stochastic gradient descent.
@@
To monitor the model performance during the training, we collected four datasets evaluating the ability of the learned embeddings to adapt to downstream tasks: 15 aryl bromides with associated reaction yields for a cross-coupling reaction3, 61 mono-substituted aryl halides with Hammett constants72, and calculated Muliken charges and NMR shielding constants. At the end of each epoch, we used the learned embeddings as features, selected the most informative features indicated by mutual information73 and the most correlated features indicated by the Pearson correlation coefficient, and trained a k-Nearest Neighbor (k-NN) to predict the reaction yields, Hammett constants, Muliken charges and NMR shielding constants. We evaluated the $r^{2}$ 74 of prediction with leave-one-out cross-validation on the four validation datasets and summed them up as a single scalar value for monitoring the training process and determining the time of early termination of the training. Training curves are shown in Figure S4.
@@
Validation with traditional reactivity descriptors To validate if the embedding learns anything about reactivity, we selected the most commonly used 500 aryl halides from the training set and supplemented them with p-, m-, and o-substituted aryl halides with common functional groups selected from a study of Hammett constants55, providing a dataset with 762 aryl halides in total. Reactivity descriptors were calculated for this set of 762 molecules. For each molecule, we computed RDKit2D descriptors75 and Morgan fingerprints59 (1024 bits, radius 2) using Therapeutic Data Commons (TDC)76. As other pre-trained representation baselines, we employed a pre-trained GIN with node-masking26 and ChemGPT37 with 4.7 M parameters, available from Deep Graph Library (DGL)77 and HuggingFace78, interfaced via molfeat79. For Figure 3B, we assessed k-Nearest Neighbor (k-NN), linear regression, L1 and L2 norm, and support vector machine (SVM)53 on each embedding, reporting the best $r^{2}$ performance. Performance validation used a 3-fold cross-validation, with mean and range of $r^{2}$ for the top model is shown in the bar plot. In Figure 3C, we trained an SVM model53 on a specific subset of the training data, using the remaining data for performance evaluation. This was repeated thrice with independent seeds, reporting the mean and range. For all machine learning models, we adopt the implementation from scikit-learn80.
@@
Downstream applications For yield prediction, we used 15 aryl bromides with experimental yields3, employing leave-one-out validation. Molecular embeddings were calculated by the ContraScope-trained GIN as previously described, and other common featurization was calculated using the same pipeline as above. For all embeddings, we selected features that are the most informative or most correlated with the prediction targets indicated by mutual information73 and Pearson correlation coefficient, and tested models including k-NN, linear regression, L1 and L2 norms, SVM, random forest, and MLP, reported the highest validation $r^{2}$. The $r^{2}$ value of the univariate model using electronegativity is from the original study3 for comparison. For regioselectivity prediction, training data comprised penta- and trifluoronitrobenzenes to predict reactivity at C-F sites in tetrafluoronitrobenzenes60. Each C–F site was encoded and labeled with its yield if reactive, or zero otherwise. During inference, all C–F sites were encoded, and the model predicted each site’s reactivity as a binary classification task. For substrate scope design, we used the same purchasable aryl bromides as Kariofillis et al. and encoded them for K-Means clustering into 15 groups, selecting a representative from each. All machine learning models were implemented using scikit-learn80.
@@
Substrate Scope Contrastive Learning: Repurposing Artificial Bias to Learn Atomic Representations
Wenhao Gao, Priyanka Raghavan, Ron Shprints and Connor W. ${}^{\ast}$
Department of Chemical Engineering, MIT, Cambridge, MA 02139
@@ 6 Code and Data Availability
All code necessary to reproduce this work can be found at https://github.com/wenhao-gao/substrate_scope_contrastive_learning/tree/main.
@@ 7 The misalignment of current pre-training methods and functionalities
Recent years have witnessed significant advancements in representation learning for molecules, accompanied by a diverse array of pre-training strategies. These strategies have emphasized various aspects, including chemical valence26, structural similarity29, and conformational information28, 30. Despite these developments, a consistent or marked enhancement in performance for downstream tasks remains elusive, as noted in Sun et al. (2022)31. A contributing factor to this challenge may be the misalignment between the objectives of these learning models and the specific requirements of their target applications, particularly in modeling molecular functionality.
Focusing on graph neural networks, early methodologies primarily targeted node or contextual prediction, as detailed in Hu et al. (2020)26. These approaches often classified atoms or groups with identical chemical valences as analogous. This technique, however, shows limitations, as evidenced in the molecules depicted in Figure S1. Here, the pre-training method erroneously identifies distinct functional groups as identical, overlooking substantial differences in their chemical functionalities. Similarly, traditional contrastive learning, which views structurally similar molecules as comparable in the embedding space29, falls short. The assumption that consistent valence bonds or structural resemblance equates to analogous molecular properties is often flawed, as significant variances may arise. While the integration of 3D conformational data offers benefits for properties reliant on specific conformations, such as those derived from quantum chemical computations28, 30, its utility diminishes in broader biological or reaction contexts. Additionally, language model pre-training that focuses on string representations like SMILES33 tends to prioritize syntax comprehension over understanding the intrinsic molecular significance.
@@ 9 Learning curves and hyper-parameter tuning
[…]
We tuned the hyper-parameter of the network architecture and the training details to maximize the summation of $r^{2}$ across these four validation tasks. The resulting GIN network has 64 hidden channels, 5 layers. An Adam optimizer was used with a momentum of 0.9, an initial learning rate of 0.00005, and a learning rate decay of 0.999. 16 triplets are sampled for each anchor molecule and a batch size of 4096 was used for training. In Equation 1, we used $\gamma$ of 4.022879258650723, $\beta$ of 0.6120957227224214, and $M$ of 1.906408074987083.

### v2

@@ 4 Discussion
Substrate scope tables from journal articles, commonly remarked upon as being being biased towards high-yielding examples, are, in fact, biased in subtler ways that correlate with chemical reactivity. By learning the patterns of what substrates were included or excluded in a publication, ContraScope extracts the underlying expert knowledge and reactivity trends used (implicitly or explicitly) to make that decision. We illustrate this both qualitatively and quantitatively, showing how our embeddings–trained only on substrate groupings in publications–align with trends in reactivity according to computed descriptors and expert intuition. To our knowledge, this is the first approach that studies publication bias as a source of information, rather than as a hindrance to modeling.
We would like to emphasize that publication bias serves as the premise and inspiration for designing the loss function without claiming or requiring that all molecules outside the reported scope tables are necessarily unreactive. Our premise that all molecules not reported in a substrate scope table are less similar to those that are reported is a simplification that holds true only on average. We acknowledge that this approach inherently introduces some false negatives during training–molecules that could undergo the reaction but are treated as negative samples. However, it is well established that machine learning models can still extract meaningful insights and achieve reasonable performance despite the presence of noisy or mislabeled data points. In our case, the model still finds signal in the patterns of substrates included in the reported scope tables that uncovers implicit reactivity trends. While excluding false negative samples through expert curation may enhance the model’s quantitative performance, addressing this challenge without conflating trends from publication bias with trends from such curation bias is complex and falls beyond the scope of this work, though some exploratory trials are included in the Supporting Information.
@@
The downstream applications showcased in this paper—such as the supervised learning of reaction yields—show the versatility and potential value of the learned ContraScope embeddings only as proofs-of-concept; the learned embedding approach, like any other, is still not a universal solution to molecular representation learning. For instance, a nearest neighbor regression analysis of yield across all training substrate scopes indicates that two molecules with similar embeddings do not necessarily exhibit similar yields, which is similar to other common features (Figure S22). Moreover, substrates may achieve dissimilar reaction yields as a result of side reactions rather than inherent dissimilarity in reactivity at the C–X motif, which would not be captured in our approach; this may partially explain the observed lack of correlation for structurally complex aryl halides in the chemistry informer library 32 (Figure S23). From a practical standpoint, our loss function combines a term for anchor-positive loss with one for anchor-negative loss in a linear manner, which can, at times, lead to instability in the training process. All of these challenges underscore that understanding and predicting chemical reactivity remains an open question, highlighting the need for further refinement of training strategies and loss functions specially designed for chemical data.
Overall, our work presents a conceptual framework for studying publication bias in a unique way through the lens of deep representation learning. We show that the resulting learned embeddings correlate with physical organic reactivity descriptors and experimental reaction performance as evidence that this information is implicit in patterns of substrate usage. This methodology can be straightforwardly extended to study trends in additional substrate classes beyond aryl halides by changing the training dataset. We are optimistic that our framework and open-source code will serve as a catalyst for further advancements in the field, fostering more functionality-aligned training strategies in molecular representation learning and the quantitative study of publication bias.
@@ 5 Methods
Substrate data for training In this study, we focus on aryl halides. We partitioned reactions recorded in the CAS Content Collection between 2010 and 2015 into substrate scope tables comprised of reactions wherein the only variable is a single aryl halide substrate. All other substrates and recorded conditions remained constant within each scope. Importantly, all reactions take place at an aryl C–X bond. Scopes with fewer than five reactions and reactions without recorded yields were excluded. This led to our final training dataset of 20,798 distinct aryl halides, covering 64,192 reactions and 6,919 substrate scopes.
Calculation of reactivity descriptors All reactivity descriptors are calculated by density-functional theory (DFT) with Gaussian 1623 via an automated descriptor generation pipeline built on top of AutoQChem24. For each unique aryl halide, up to 20 conformers were generated using RDKit’s ETKDG algorithm33 and optimized using the MMFF94 force field34. To reduce computational overhead when using DFT, the lowest-energy conformer for each molecule was then selected via GFN2-xTB35. Any conformer for which any energy calculation did not converge was discarded. The lowest-energy conformer for each aryl halide then underwent geometry optimization and frequency calculations in Gaussian 16 using the B3LYP functional36, 37 with the 6-31G*38 basis set. Atoms with atomic number $>35$ use the LANL2DZ39 basis set instead. This led to the generation of 25 molecule-level descriptors per molecule (energies, energy corrections, dipole moment, HOMO/LUMO energies, electronegativity, etc.) and 19 atom-level descriptors per atom per molecule (buried volume, partial charges, NMR shielding constants, etc.), and the relevant atom-level descriptors for the reactive C–X bond were extracted. The full set of descriptors is listed in Figures S12 and S13.
Featurization and network architecture We adopt a graph representation of molecules and the graph isomorphism network (GIN)18 for modeling them. GIN is one kind of graph neural network that updates the representation of each atom over multiple iterations as follows:
@@
where $h_{v}^{(k)}$ is the atom representation of atom $v$ at $k$-th layer, $\mathcal{N}(v)$ is a set of atoms that are covalently bonded to atom $v$. $\epsilon$ is a learnable scalar and MLP represents a multi-layer preceptrons model. Utilizing this atom-wise message-passing schema, GIN is able to propagate and aggregate information through network layers to embed atoms in a manner that respects graph isomorphism. After hyper-parameter tuning, the network has a hidden dimension of 64, and message passing layer of 5. For the initial representation of molecular graphs, we use minimal featurizations19 that include atomic numbers, total degrees, formal charges, chiral tags, number of hydrogens, hybridization types, aromaticity, mass, and bond types.
Substrate scope contrastive learning The training of substrate scope contrastive learning involves triplet sampling, computing loss function depicted in Eq. 1, and back-propagation. Within an epoch, each aryl halide serves as the anchor molecule once and 16 triplets are sampled randomly for each anchor. We exclude all cases with identical molecules sampled in one triplet. Based on empirical performance, we adopt a distance metric called signal-noise ratio distance (SNR) 40, defined as follows:
@@
where ${f}_{i},{f}_{j}$ is the embedding of aryl halide $m_{i},m_{j}$. We take the representation of an aryl halide’s reactivity, $f_{i}$, to be the atom-level feature of the carbon atom in the reactive C–X bond from the GIN. The Adam41 optimization algorithm is used for stochastic gradient descent.
@@
To monitor the model performance during the training, we collected four datasets evaluating the ability of the learned embeddings to adapt to downstream tasks: 15 aryl bromides with associated reaction yields for a cross-coupling reaction8, 61 mono-substituted aryl halides with Hammett constants42, and calculated Muliken charges and NMR shielding constants. At the end of each epoch, we used the learned embeddings as features, selected the most informative features indicated by mutual information43 and the most correlated features indicated by the Pearson correlation coefficient, and trained a k-Nearest Neighbor (k-NN) to predict the reaction yields, Hammett constants, Muliken charges and NMR shielding constants. We evaluated the $r^{2}$ 44 of prediction with leave-one-out cross-validation on the four validation datasets and summed them up as a single scalar value for monitoring the training process and determining the time of early termination of the training. Training curves are shown in Figure S8.
@@
Regression analysis with traditional reactivity descriptors To validate if the embedding learns anything about reactivity, we selected the most commonly used 500 aryl halides from the training set and supplemented them with p-, m-, and o-substituted aryl halides with common functional groups selected from a study of Hammett constants22, providing a dataset with 762 aryl halides in total. Reactivity descriptors were calculated for this set of 762 molecules. For each molecule, we computed RDKit2D descriptors45 and Morgan fingerprints27 (1024 bits, radius 2) using Therapeutic Data Commons (TDC)46. As for pre-trained representation baselines, we adopted AttrMask, 3DInfoMax, and ChemGPT. Among them, AttrMaskis a pre-trained GIN with node-masking28 and ChemGPT30 is a chemical language model with 4.7 M parameters, both available from Deep Graph Library (DGL)47 and HuggingFace48, interfaced via molfeat49. 3DInfoMax29 is a graph neural network model pre-trained to incorporate information from three-dimensional molecular conformations. In our work, we adopt the pre-trained parameters provided by the original repository to ensure consistency. For Figure 3B, we assessed k-Nearest Neighbor (k-NN), linear regression, L1 and L2 norm, and support vector machine (SVM)25 on each embedding, reporting the best $r^{2}$ performance. Performance validation used a 3-fold cross-validation, with mean and range of $r^{2}$ for the top model is shown in the bar plot. In Figure 3C, we trained an SVM model25 on a specific subset of the training data, using the remaining data for performance evaluation. This was repeated thrice with independent seeds, reporting the mean and range. For all machine learning models, we adopt the implementation from scikit-learn50.
@@
Downstream applications For yield prediction, we used 15 aryl bromides with experimental yields8, employing leave-one-out validation. Molecular embeddings were calculated by the ContraScope-trained GIN as previously described, and other common featurization was calculated using the same pipeline as above. For all embeddings, we selected features that are the most informative or most correlated with the prediction targets indicated by mutual information43 and Pearson correlation coefficient, and tested models including k-NN, linear regression, L1 and L2 norms, SVM, random forest, and MLP, reported the highest validation $r^{2}$. The $r^{2}$ value of the univariate model using electronegativity is from the original study8 for comparison. For regioselectivity prediction, training data comprised penta- and trifluoronitrobenzenes to predict reactivity at C-F sites in tetrafluoronitrobenzenes31. Each C–F site was encoded and labeled with its yield if reactive, or zero otherwise. During inference, all C–F sites were encoded, and the model predicted each site’s reactivity as a binary classification task. For substrate scope design, we used the same purchasable aryl bromides as Kariofillis et al. and encoded them for K-Means clustering into 15 groups, selecting a representative from each. All machine learning models were implemented using scikit-learn50.
6 Supporting Information
The Supporting Information includes links to the code and validation data, a discussion on the differences between the introduced method and existing pre-training approaches, additional statistics on the training dataset, and further experimental and computational results.
@@ 6 Supporting Information
Revealing the Relationship Between Publication Bias and Chemical Reactivity with Contrastive Learning
Wenhao Gao, Priyanka Raghavan, Ron Shprints, and Connor W. Coley,∗,†,‡
$\dagger$ Department of Chemical Engineering, Massachusetts Institute of Technology, Cambridge, MA 02139, United States
$\ddagger$ Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA 02139, United States
@@ 7 Code and Data Availability
All code, pretrained model weights, and validation data necessary to reproduce the major results of this work can be found at https://github.com/wenhao-gao/substrate_scope_contrastive_learning/tree/main. We are unable to share the full dataset from the CAS Content Collection that contains the full substrate scope information used for contrastive learning.
@@ 8 The misalignment of current pre-training methods and functionalities
Recent years have witnessed significant advancements in representation learning for molecules, accompanied by a diverse array of pre-training strategies. These strategies have emphasized various aspects, including chemical valence28, structural similarity51, and conformational information52, 29. Despite these developments, a consistent or marked enhancement in performance for downstream tasks remains elusive, as noted in Sun et al. (2022)53. A contributing factor to this challenge may be the misalignment between the objectives of these learning models and the specific requirements of their target applications, particularly in modeling molecular functionality.
Focusing on graph neural networks, early methodologies primarily targeted node or contextual prediction, as detailed in Hu et al. (2020)28. These approaches often classified atoms or groups with identical chemical valences as analogous. This technique, however, shows limitations, as evidenced in the molecules depicted in Figure S1. Here, the pre-training method erroneously identifies distinct functional groups as identical, overlooking substantial differences in their chemical functionalities. Similarly, traditional contrastive learning, which views structurally similar molecules as comparable in the embedding space51, falls short. The assumption that consistent valence bonds or structural resemblance equates to analogous molecular properties is often flawed, as significant variances may arise. While the integration of 3D conformational data offers benefits for properties reliant on specific conformations, such as those derived from quantum chemical computations52, 29, its utility diminishes in broader biological or reaction contexts. Additionally, language model pre-training that focuses on string representations like SMILES54 tends to prioritize syntax comprehension over understanding the intrinsic molecular significance.
@@ 9 Statistics of the dataset
Figure S4: Bar plot showing the distribution of aromatic ring sizes in the training dataset.
Figure S5: Bar plot illustrating the distribution of different halide types in the training dataset.
Figure S6: The bar plot displays the 30 reaction types with the number of occurrences in our training data, illustrating their relative prevalence. Each reaction type is labeled on the x-axis, with the number of occurrences represented on the y-axis.
Figure S7: Examples of five-membered aromatic rings present in our training dataset.
@@ 10 Learning curves and hyper-parameter tuning
The GIN network was trained with our substrate scope contrastive loss, as described in Equation 1. The value of the total loss, anchor-positive term, and anchor-negative term are depicted in Figure S8(a-c). The cumulative gradient of the total loss across all trainable parameters is illustrated in Figure S8(d). As described in the Method section, under the Substrate scope contrastive learning subsection, we constructed four validation tasks to monitor the training process. This monitoring involved evaluating the network’s ability to predict reaction yields, Hammett constants, Mulliken charges, and NMR shielding constants. The summation of the coefficient of determination ($r^{2}$) values across these four validation tasks, as well as the Pearson correlation coefficient between the predicted targets and their most correlated features, are presented in Figure S8(e-f). Overall, we can see the $r^{2}$ and the Pearson correlation coefficient plateau after 50 epochs and oscillate after that. We terminated the training at the 56th epoch, which reached the highest aggregate $r^{2}$ value.
We tuned the hyper-parameter of the network architecture and the training details to maximize the summation of $r^{2}$ across these four validation tasks. The resulting GIN network has 64 hidden channels and 5 layers. An Adam optimizer was used with a momentum of 0.9, an initial learning rate of 0.00005, and a learning rate decay of 0.999. 16 triplets were sampled for each anchor molecule, and a batch size of 4096 was used for training. In Equation 1, we used $\gamma$ of 4.023, $\beta$ of 0.612, and $M$ of 1.906.
@@
We show the correlation between the learned embeddings and the Hammett constants in Figures S8–S10. It is important to note that, even for molecules seen during pretraining, the model does not have access to the Hammett constant values, and all metrics reported are for the “test set” in the conventional supervised learning setting. This demonstrates that the model has learned transferable patterns reflecting reactivity through substrate scope contrastive learning.
We also want to highlight that 42 of the 61 molecules with labeled Hammett constants were seen during pretraining. This is because our pretraining dataset includes all reactions involving aryl halides with at least five substrates, which means that most commonly used molecules were already encountered. Consequently, there may be a distribution shift between molecules seen during pretraining and those not seen.
Figure S9: Results of predicting Hammett constants from learned embeddings.
Figure S10: Results of predicting Hammett constants from learned embeddings. This figure shows a subset of data where the aryl halides are seen during pre-training.
Figure S11: Results of predicting Hammett constants from learned embeddings. This figure shows a subset of data where the aryl halides are not seen during pre-training.